OpenStudy (anonymous):

Even though the Moon is much smaller than the Earth, it still exerts a pull of gravity on our planet. Explain what effect the Moon’s pull has on the Earth, and why this occurs.

OpenStudy (anonymous):

Please help?!

OpenStudy (anonymous):

The Moon's gravity affects everything on Earth. It calls not only the waters but the very Earth. Even though the Moon is so much smaller than the Sun, it is much closer to us. So it exerts over twice the gravitational force on the Earth than does the Sun. And these lunar influences are a vital factor in the Earth's ability to support life. Science lists 4 Moon-tides. I include one that science prefers to overlook, no matter how unscientifically: the 5th Moon-tide. . . the tidal pull on Spirit.

OpenStudy (anonymous):

Thank you!!

OpenStudy (stormfire1):

@lameer1: The text you cut and pasted is misleading because it says the moon's gravitational force is twice that of the sun which is completley wrong. It should have said "tidal forces". So the original poster understands the difference: The general formula for gravitational force (in Newtons) is: \[F=G\frac{M_1M2}{r^2}\]where F is the force, G is the gravitational constant (in N(m/kg)^2), M1 and M2 are the relative masses (in kg) and r is the average center-to-center distance between the masses (in km). From those formulas you get: \[F_{grav_{MOON-EARTH}}=6.67x10^{-11}N(m/kg)^2\frac{(7.36 x 10^{22}kg)(5.97x10^{24} kg)}{(3.84x10^5km)^2}=1.99x10^{26}N\]\[F_{grav_{SUN-EARTH}}=6.67x10^{-11}N(m/kg)^2\frac{(1.99 x 10^{30}kg)(5.97x10^{24} kg)}{(1.5x10^8km)^2}=3.52x10^{28}N\]Divide out the forces and you'll see that the Sun exterts ~177 times the gravitational force on the earth than the moon does. The tidal forces however are calculated differently:\[F_{tidal}=2G\frac{M_1 M_2R}{r^3}\]The fact that the distance is cubed here explains why the moon exerts more *tidal forces* on the earth than the sun...which is what that original statement probably meant to say.

OpenStudy (stormfire1):

@alexis: check out this video for a decent explanation of the moon and tides http://www.youtube.com/watch?v=aN2RM5wa1ek